# New versions of Frobenius and integral closure of ideals

### v2

@@ 3.2. The $\mathfrak{a}^{t}$-tight closures
Throughout this subsection, we assume $R$ is of prime characteristic $p$. Recall the following definitions from [HY03] and [Vra08].
@@
Let $I,\mathfrak{a}\subset R$ be ideals and let $t\in\mathbb{R}_{>0}$. Then, the Hara-Yoshida $\mathfrak{a}^{t}$-tight closure of $I$, denoted $I^{*\mathfrak{a}^{t}},$ is defined by
$I^{*\mathfrak{a}^{t}}=\left\{x\in R\left|\text{ for some }c\in R^{\circ}\text{and all }e\gg 0,\,cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{\left[p^{e}\right]}\right.\right\}.$
@@
${}^{\mathfrak{a}^{t}}I^{*}=\left\{x\in R\left|\text{ for some }c\in R^{\circ}\text{ and all }e\gg 0,\,cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset\mathfrak{a}^{\lceil tp^{e}\rceil}I^{\left[p^{e}\right]}\right.\right\}.$
@@
Let $I,\mathfrak{a}\subset R$ be ideals and let $t\in\mathbb{Q}_{>0}$. Then $I^{*\mathfrak{a}^{t}}\subset(I^{*}:\mathfrak{a}_{t})$.
@@
The following result shows that, like tight closure, the computation of $\mathfrak{a}^{t}$-tight closure reduces to computation in domains via modding out by minimal primes. Note that while we mention the result for the $t=1$ case, it holds similarly for $t\neq 1$ as well.
@@
Let $\mathfrak{a},I\subset R$ be ideals with $\mathfrak{a}\cap R^{\circ}\neq\varnothing.$ Then, $x\in$ ${}^{\mathfrak{a}}I^{*}$ ($x\in I^{*\mathfrak{a}}$) if and only if, for each minimal prime $\mathfrak{p}$ of $R$, we have $x+\mathfrak{p}$ in ((${}^{\mathfrak{b}}((I+\mathfrak{p})/\mathfrak{p})^{*}$ (or $((I+\mathfrak{p})/\mathfrak{p})^{*\mathfrak{b}}$), where $\mathfrak{b}$ denotes the image of $\mathfrak{a}$ in $R/\mathfrak{p}.$
@@
Let $x\in$ ${}^{\mathfrak{a}}I^{*}.$ By definition, there exists $c\in R^{\circ}$ such that $cx^{p^{e}}\mathfrak{a}^{p^{e}}\subset\mathfrak{a}^{p^{e}}I^{[p^{e}]}$ for all large $e.$ Thus, for any minimal prime $\mathfrak{p}$ of $R,$ image of $c$ in $R/\mathfrak{p}$ is in $(R/\mathfrak{p})^{\circ}$ and $cx^{p^{e}}\mathfrak{a}^{p^{e}}+\mathfrak{p}\subset\mathfrak{a}^{p^{e}}I^{[p^{e}]}+\mathfrak{p}$ for all large $e$ implying that the image of $x$ in $R/\mathfrak{p}$ is in ((${}^{\mathfrak{b}}((I+\mathfrak{p})/\mathfrak{p})^{*}.$
Conversely, assume that for each minimal prime $\mathfrak{p}$ of $R,$ image of $x$ in $R/\mathfrak{p}$ is in ((${}^{\mathfrak{b}}((I+\mathfrak{p})/\mathfrak{p})^{*}.$ Let $c_{\mathfrak{p}}\in R\setminus\mathfrak{p}$ such that $c_{\mathfrak{p}}x^{p^{e}}\mathfrak{a}^{p^{e}}\subset c_{\mathfrak{p}}x^{p^{e}}\mathfrak{a}^{p^{e}}+\mathfrak{p}\subset\mathfrak{a}^{p^{e}}I^{[p^{e}]}+\mathfrak{p}$ for all $e$ large. By Prime Avoidance, there exists a non-zero element $d_{\mathfrak{p}}\in R$ which is not in $\mathfrak{p}$ but is in every other minimal prime of $R.$ Then $d_{\mathfrak{p}}c_{\mathfrak{p}}x^{p^{e}}\mathfrak{a}^{p^{e}}\subset\mathfrak{a}^{p^{e}}I^{[p^{e}]}+\sqrt{0},$ for all $e$ large. Set $c^{\prime}=\sum_{\mathfrak{p}}d_{\mathfrak{p}}c_{\mathfrak{p}}.$ Then $c^{\prime}\in R^{\circ}$ and for all $e$ large,
$c^{\prime}x^{p^{e}}\mathfrak{a}^{p^{e}}\subset\mathfrak{a}^{p^{e}}I^{[p^{e}]}+\sqrt{0}.$
Since $R$ is a Noetherian ring, there exists $e^{\prime}$ such that $(\sqrt{0})^{[p^{e^{\prime}}]}=(0).$ Thus, for all $e$ large, we have $c^{\prime\prime}x^{p^{e+e^{\prime}}}(\mathfrak{a}^{p^{e}})^{[p^{e^{\prime}}]}\subset(\mathfrak{a}^{p^{e}})^{[p^{e^{\prime}}]}I^{[p^{e+e^{\prime}}]}$. From [Tay21, Lemma 2.1], we know that $\mathfrak{a}^{p^{e+e^{\prime}}}\mathfrak{a}^{(\mu-1)p^{e^{\prime}}}\subset(\mathfrak{a}^{[p^{e^{\prime}}]})^{p^{e}},$ where $\mu$ denotes the minimal number of generators of $\mathfrak{a}.$ Hence, for any $b\in\mathfrak{a}^{(\mu-1)p^{e^{\prime}}}\cap R^{\circ},$ and for all $e$ large,
$bc^{\prime\prime}x^{p^{e+e^{\prime}}}\mathfrak{a}^{p^{e+e^{\prime}}}\subset c^{\prime\prime}x^{p^{e+e^{\prime}}}(\mathfrak{a}^{[p^{e^{\prime}}]})^{p^{e}}\subset(\mathfrak{a}^{p^{e}})^{[p^{e^{\prime}}]}I^{[p^{e+e^{\prime}}]}\subset\mathfrak{a}^{p^{e+e^{\prime}}}I^{[p^{e+e^{\prime}}]}.$
As $bc^{\prime\prime}\in R^{\circ},$ we get $x\in$ ${}^{\mathfrak{a}}I^{*}$. The proof for the ideal $I^{*\mathfrak{a}}$ follows similarly. ∎
@@ 3.3. The $\mathfrak{a}^{t}$-Frobenius closures
Throughout this subsection, we assume $R$ is of prime characteristic $p$.
@@
Let $I,\mathfrak{a}\subset R$ be ideals and let $t\in\mathbb{R}_{>0}$. Then, the Hara-Yoshida $\mathfrak{a}^{t}$-Frobenius closure of $I$, denoted $I^{F\mathfrak{a}^{t}},$ is defined by
$I^{F\mathfrak{a}^{t}}=\left\{x\in R\left|\text{ for all }e\gg 0,\,x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{\left[p^{e}\right]}\right.\right\}.$
@@
${}^{\mathfrak{a}^{t}}I^{F}=\left\{x\in R\left|\text{ for all }e\gg 0,\,x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset\mathfrak{a}^{\lceil tp^{e}\rceil}I^{\left[p^{e}\right]}\right.\right\}.$
@@
Let $I,\mathfrak{a}\subset R$ be ideals and let $t\in\mathbb{R}_{>0}$. Then, we have the following.
@@
    (${}^{\mathfrak{a}^{t}}({}^{\mathfrak{a}^{t}}I^{F})^{F}={}^{\mathfrak{a}^{t}}I^{F}.$
@@
    Notice that $x\in I^{F\mathfrak{a}^{t}}$ if and only if $x^{p^{e}}\in I^{\left[p^{e}\right]}:\mathfrak{a}^{\lceil tp^{e}\rceil}$ for all $e\gg 0$, from whence it is obvious that $I^{F\mathfrak{a}^{t}}$ is an ideal of $R$. Similarly, $x\in{}^{\mathfrak{a}^{t}}I^{F}$ if and only if $x^{p^{e}}\in I^{\left[p^{e}\right]}\mathfrak{a}^{\lceil tp^{e}\rceil}:\mathfrak{a}^{\lceil tp^{e}\rceil}$, so ${}^{\mathfrak{a}^{t}}I^{F}$ is also an ideal of $R$.
@@
    From part b), it follows that ${}^{\mathfrak{a}^{t}}I^{F}\subset{}^{\mathfrak{a}^{t}}({}^{\mathfrak{a}^{t}}I^{F})^{F}.$ Now, let $x\in{}^{\mathfrak{a}^{t}}({}^{\mathfrak{a}^{t}}I^{F})^{F}.$ Then for all $e\gg 0,$ $x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset\mathfrak{a}^{\lceil tp^{e}\rceil}({}^{\mathfrak{a}^{t}}I^{F})^{[p^{e}]}\subset\mathfrak{a}^{\lceil tp^{e}\rceil}I^{[p^{e}]},$ where the last containment holds by checking the generators of ${}^{\mathfrak{a}^{t}}I^{F}$ and using that ${}^{\mathfrak{a}^{t}}I^{F}$ is a finitely generated $R$-module. ∎
@@
Unlike $\mathfrak{a}^{t}$-integral closure and $\mathfrak{a}^{t}$-tight closure, $\mathfrak{a}^{t}$-Frobenius closure can’t be computed using the corresponding closure ideal after going modulo the minimal primes. We use 2.3 to check that 3.11 may not hold true for ${}^{\mathfrak{a}^{t}}I^{F}$ and $I^{F\mathfrak{a}^{t}}$.
@@
Let $R=\mathbb{F}_{3}[X,Y,Z]/(Y^{6}-X^{3},X^{2}-Z^{2})$ and let $x,y,$ and $z$ denote the images of $X,Y,$ and $Z$ respectively in $R.$ Let $I=(x)$ and $\mathfrak{a}=(x,y^{2})$. We claim that $z\notin$ ${}^{\mathfrak{a}}I^{F}.$ Indeed, if $z\in$ ${}^{\mathfrak{a}}I^{F},$ then $z^{3^{e}}(x,y^{2})^{3^{e}}\subset\mathfrak{a}^{3^{e}}I^{[3^{e}]}$ for some $e.$ In particular, $z^{3^{e}}x^{3^{e}}\in\mathfrak{a}^{3^{e}}I^{[3^{e}]}=(x,y^{2})^{3^{e}}(x^{3^{e}})$ for some $e.$ Write $13^{e}=2k+$ 1 for some $k.$ Then
$z^{3^{e}}x^{3^{e}}=z^{2k+1}x^{3^{e}}=z(z^{2})^{k}x^{3^{e}}=zx^{2k}x^{3^{e}}=zx^{3^{e}-1}x^{3^{e}}.$
Since $x$ is a nonzerodivisor, we get $z\in(x,y^{2})^{3^{e}}:x^{3^{e}-1}=(x,y^{2})$ which is not true. Thus, the claim holds. Note that $\operatorname{Min}R=\{(x-z,y^{2}-z),(x+z,y^{2}+z)\}$ and so modulo either minimal prime, the image of $z$ is the same as the image of $x.$ Thus, $z+\mathfrak{p}\in((x)+\mathfrak{p})/\mathfrak{p}\subset$ ((${}^{\mathfrak{a}}((x)+\mathfrak{p})^{F}/\mathfrak{p}$, but $z\not\in$ (${}^{\mathfrak{a}}(x)^{F}$.
Now let $I=(x+y^{2})$ and $\mathfrak{a}=(x,y^{2}).$ Then one can check that modulo any minimal prime $\mathfrak{p},$ the element $(1+\mathfrak{p})/\mathfrak{p}$ is in $((I+\mathfrak{p})/\mathfrak{p})^{F\mathfrak{a}}$ but not in $I^{F\mathfrak{a}}.$ In particular, it is enough to show that $(\mathfrak{a}^{3^{e}}+\mathfrak{p})/\mathfrak{p}\subset(I^{[3^{e}]}+\mathfrak{p})/\mathfrak{p}$ for every minimal prime $\mathfrak{p}$ and for all $e$ large but $\mathfrak{a}^{3^{e}}\not\subset I^{[3^{e}]}$ for any $e.$ Indeed, modulo any minimal prime $\mathfrak{p},$
@@
As for the $\mathfrak{a}^{t}$-tight and integral closures, if $\mathfrak{a}=R$, then $I^{FR^{t}}=\,^{R^{t}}I^{F}=I^{F}$.
@@
    Let $t>\mu(\mathfrak{a})$ a real number, where $\mu(\mathfrak{a})$ denotes the minimal number of generators of $\mathfrak{a}$. Then
@@
    Thus, $(I^{F(f)})^{F(f)}=(I^{F}:(f)^{2})$ which is strictly larger than $I^{F(f)}=I^{F}:(f)$ under the given assumptions.
@@
3.4. Comparing the new closures
In this subsection, we compare the different adjusted closures with one another. All statements which involve the adjusted integral closures only are characteristic independent, but whenever the adjusted tight or Frobenius closures are mentioned, we tacitly assume the ambient ring is of prime characteristic $p>0$. Furthermore, in any statement where we invoke the adjusted integral closures, we assume $I\cap R^{\circ}$ and $\mathfrak{a}\cap R^{\circ}$ are nonempty to ensure the adjusted closures are ideals.
By taking $c=1$ in the definition of the (adjusted) $\mathfrak{a}^{t}$-tight closures, we can easily see $I^{F\mathfrak{a}^{t}}\subset I^{*\mathfrak{a}^{t}}$ and $\,{}^{\mathfrak{a}^{t}}I^{F}\subset\,^{\mathfrak{a}^{t}}I^{*}$. Similarly, the adjusted $\mathfrak{a}^{t}$-tight closures are inside the corresponding adjusted $\mathfrak{a}^{t}$-integral closure. In general, we have the following inclusions of ideal closures, fixing $I$, $\mathfrak{a}$, and $t>0$, and assuming $I\cap R^{\circ}$ and $\mathfrak{a}\cap R^{\circ}$ are nonempty.

### v3

@@ 3.1. The $\mathfrak{a}^{t}$-integral closures
3.1.1. Valuative Theory of $\mathfrak{a}^{t}$-Integral Closure
We now explore the connections between our new integral closure operation and valuation theory. As for ordinary integral closure, we will use the Rees valuations of an ideal $I$ to understand $\overline{I}^{\mathfrak{a}^{t}}$. For the sake of concision, we will avoid a complete review of Rees valuations here, and direct the reader to [SH06, Chapter 10] for a thorough treatment. However, we will record some of their properties below.
Each ideal $I$ in a noetherian ring $R$ has a finite set of valuations $\mathcal{RV}(I)$ with the following properties ([Ree56]). [^2] Following [SH06, page 192], we slightly abuse the term “valuation” by allowing the functions to be defined on non-domains and to take the value $\infty$ on certain minimal primes.
  1. (1)
    Each $\nu\in\mathcal{RV}(I)$ is a valuation, i.e. a function $\nu:R\setminus\{0\}\to\mathbb{Z}_{\geq 0}\cup\{\infty\}$ such that for each $0\neq x,y\in R$, $\nu(xy)=\nu(x)+\nu(y)$ and $\nu(x+y)\geq\min\{\nu(x),\nu(y)\}$. For a nonzero ideal $J\subset R$ and Rees valuation $\nu\in\mathcal{RV}(I)$, we define $\nu(J)=\min\{\nu(x)\mid x\in J\}$.
  2. (2)
    For $x\in R$, $x\in\overline{I^{n}}$ if and only if $\nu(x)\geq n\nu(I)$ for all $\nu\in\mathcal{RV}(I)$ (see [SH06, Remark 10.1.4]).
  3. (3)
    $J\subset\overline{I^{n}}$ if and only if $\nu(J)\geq n\nu(I)$.
Theorem 3.9.
Let $R$ be a ring, $I,\mathfrak{a}\subset R$ ideals of positive height, and $t\in\mathbb{R}_{>0}$. Let $\mathcal{RV}(I)$ be the set of Rees valuations of $I$. An element $x\in R$ is in $\overline{I}^{\mathfrak{a}^{t}}$ if and only if for all $\nu\in\mathcal{RV}(I)$, $\nu(x)\geq\nu(I)-t\nu(\mathfrak{a})$.
Proof.
Suppose $x\in\overline{I}^{\mathfrak{a}^{t}}$, and let $\nu\in\mathcal{RV}(I)$. There exists $c\in R^{\circ}$ such that for $n\gg 0$, $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}\subset\overline{I^{n}}$. Applying $\nu$ to both sides and dividing by $n$, we have
$\frac{\nu(c)}{n}+\nu(x)+\frac{\lceil tn\rceil}{n}\nu(\mathfrak{a})\geq\nu(I).$
Then, taking the limit as $n\to\infty$ gives $\nu(x)+t\nu(\mathfrak{a})\geq\nu(I)$.
Now suppose that $\nu(x)\geq\nu(I)-t\nu(\mathfrak{a})$ for all $\nu\in\mathcal{RV}(I)$. Choose $s\leq t$ rational such that $\nu(x)\geq\nu(I)-s\nu(\mathfrak{a})$ for all $\nu\in\mathcal{RV}(I)$. Let $m\in\mathbb{N}$ such that $sm\in\mathbb{N}$. We have that
$\nu(x^{m}\mathfrak{a}^{sm})=m\nu(x)+sm\nu(\mathfrak{a})=m(\nu(x)+s\nu(\mathfrak{a}))\geq m\nu(I).$
Therefore $x^{m}\mathfrak{a}^{sm}\subset\overline{I^{m}}$. Thus there exists $c^{\prime}\in R^{\circ}$ such that for all $n\in\mathbb{N}$, $c^{\prime}\left(x^{m}\mathfrak{a}^{sm}\right)^{n}\subset I^{mn}$.
Let $c^{\prime\prime}\in I^{m-1}\cap R^{\circ}$ and $c=c^{\prime}c^{\prime\prime}\in R^{\circ}$. Now, for any $N\gg 0$, let $n\in\mathbb{N}$ such that $mn\leq N<m(n+1)$, and notice
$cx^{N}\mathfrak{a}^{\lceil sN\rceil}\subset c^{\prime}c^{\prime\prime}x^{mn}\mathfrak{a}^{tmn}\subset c^{\prime\prime}I^{mn}\subset I^{mn+m-1}\subset I^{N}.$
Therefore $x\in\overline{I}^{\mathfrak{a}^{s}}\subset\overline{I}^{\mathfrak{a}^{t}}$ by $3.$ ∎
The valuative criterion allows us to show that containment in the Hara-Yoshida integral closure of an ideal $I$ can be checked using infinitely many powers rather than all sufficiently large powers, which is also true for the ordinary integral closure (see [SH06, Cor. 6.8.12]).
Corollary 3.10.
Let $R$ be a ring, $I,\mathfrak{a}\subset R$ ideals of positive height, $t\in\mathbb{R}_{>0}$, and $x\in R$. If there exists $c\in R^{\circ}$ such that $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$ for infinitely many $n\in\mathbb{N}$, then $x\in\overline{I}^{\mathfrak{a}^{t}}$.
Proof.
Let $\nu$ be a Rees valuation of $I$. For infinitely many $n\in\mathbb{N}$, $cx^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{n}$, and therefore $\nu(c)+n\nu(x)+\lceil tn\rceil\nu(\mathfrak{a})\geq n\nu(I)$, implying $\nu(x)\geq\nu(I)-\dfrac{\lceil tn\rceil}{n}\nu(\mathfrak{a})-\dfrac{\nu(c)}{n}$. Therefore
$\nu(x)\geq\liminf_{n\to\infty}\left(\nu(I)-\dfrac{\lceil tn\rceil}{n}\nu(\mathfrak{a})-\dfrac{\nu(c)}{n}\right)=\nu(I)-t\nu(\mathfrak{a}).$
Since this holds for all Rees valuations $\nu$ of $I$, $x\in\overline{I}^{\mathfrak{a}^{t}}$ by Theorem 3.9. ∎
Theorem 3.9 is also related to a description of the $\mathfrak{a}^{t}$-integral closure of a monomial ideal in an affine semigroup ring, see Theorem 6.8.
@@ 3.2. The $\mathfrak{a}^{t}$-tight closures
Throughout this subsection, we assume $R$ is of prime characteristic $p>0$. Recall the following definitions from [HY03] and [Vra08].
@@
Let $R$ be a ring, $I,\mathfrak{a}\subset R$ be ideals, and $t\in\mathbb{R}_{>0}$. Then, the Hara-Yoshida $\mathfrak{a}^{t}$-tight closure of $I$, denoted $I^{*\mathfrak{a}^{t}},$ is defined by
[…]
@@
[…]
Note that the Hara-Yoshida $\mathfrak{a}^{t}$-tight closure is only a pre-closure ([Vra08, Observation 2.3(2)]) whereas the Vraciu $\mathfrak{a}^{t}$-tight closure is an ideal closure operation ([Vra08, Observation 2.3(4)]).
@@
Let $R$ be a ring, $I,\mathfrak{a}\subset R$ be ideals and $t\in\mathbb{Q}_{>0}$. Then $I^{*\mathfrak{a}^{t}}\subset(I^{*}:\mathfrak{a}_{t})$.
@@
The following result shows that, like tight closure, the computation of $\mathfrak{a}^{t}$-tight closure reduces to computation in domains via modding out by minimal primes.
@@
[…]
@@
[…]
$c^{\prime}x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset\mathfrak{a}^{\lceil tp^{e}\rceil}I^{[p^{e}]}+\sqrt{0}.$
If $R$ is reduced, then we’re done since $\sqrt{0}=(0)$. So, assume that $\mathfrak{a}$ is of positive height. Since $R$ is a noetherian ring, there exists $e^{\prime}$ such that $(\sqrt{0})^{[p^{e^{\prime}}]}=(0).$ Thus, for all $e$ large, we have $(c^{\prime})^{p^{e^{\prime}}}x^{p^{e+e^{\prime}}}(\mathfrak{a}^{\lceil tp^{e}\rceil})^{[p^{e^{\prime}}]}\subset(\mathfrak{a}^{\lceil tp^{e}\rceil})^{[p^{e^{\prime}}]}I^{[p^{e+e^{\prime}}]}$. From [Tay21, Lemma 2.1], we know that $\mathfrak{a}^{\lceil tp^{e+e^{\prime}}+\mu(\mathfrak{a})p^{e^{\prime}}\rceil}\subset(\mathfrak{a}^{\lceil tp^{e}\rceil})^{[p^{e^{\prime}}]}$. So, let $b\in\mathfrak{a}^{\mu(\mathfrak{a})p^{e^{\prime}}}\cap R^{\circ}$. For all $e$ large,
$b(c^{\prime})^{p^{e^{\prime}}}x^{p^{e+e^{\prime}}}\mathfrak{a}^{\lceil tp^{e+e^{\prime}}\rceil}\subset(c^{\prime})^{p^{e^{\prime}}}x^{p^{e+e^{\prime}}}(\mathfrak{a}^{\lceil tp^{e}\rceil})^{[p^{e^{\prime}}]}\subset(\mathfrak{a}^{\lceil tp^{e}\rceil})^{[p^{e^{\prime}}]}I^{[p^{e+e^{\prime}}]}\subset\mathfrak{a}^{\lceil tp^{e+e^{\prime}}\rceil}I^{[p^{e+e^{\prime}}]}.$
Since $b(c^{\prime})^{p^{e^{\prime}}}\in R^{\circ}$, this shows $x\in$ $^{\mathfrak{a}^{t}}I^{*}$. The proof for the ideal $I^{*\mathfrak{a}^{t}}$ follows similarly. ∎
@@ 3.3. The $\mathfrak{a}^{t}$-Frobenius closures
Throughout this subsection, we assume $R$ is of prime characteristic $p>0$.
@@
Let $R$ be a ring, $I,\mathfrak{a}\subset R$ be ideals, and $t\in\mathbb{R}_{>0}$. Then, the Hara-Yoshida $\mathfrak{a}^{t}$-Frobenius closure of $I$, denoted $I^{F\mathfrak{a}^{t}},$ is defined by
[…]
@@
[…]
As before, we note that the Hara-Yoshida $\mathfrak{a}^{t}$-Frobenius closure is a pre-closure (Proposition 3.15) and not an ideal closure operation (Proposition 3.18(a)), whereas the Vraciu $\mathfrak{a}^{t}$-Frobenius closure is an ideal closure operation (Proposition 3.15).
@@
Let $R$ be a ring, $I,\mathfrak{a}\subset R$ be ideals, and $t\in\mathbb{R}_{>0}$. Then, we have the following.
@@
    $^{\mathfrak{a}^{t}}({}^{\mathfrak{a}^{t}}I^{F})^{F}={}^{\mathfrak{a}^{t}}I^{F}.$
@@
    Notice that $x\in I^{F\mathfrak{a}^{t}}$ if and only if $x^{p^{e}}\in I^{\left[p^{e}\right]}:\mathfrak{a}^{\lceil tp^{e}\rceil}$ for all $e\gg 0$, from whence it is obvious that $I^{F\mathfrak{a}^{t}}$ is an ideal of $R$. Similarly, $x\in{}^{\mathfrak{a}^{t}}I^{F}$ if and only if $x^{p^{e}}\in I^{\left[p^{e}\right]}\mathfrak{a}^{\lceil tp^{e}\rceil}:\mathfrak{a}^{\lceil tp^{e}\rceil}$ for all $e\gg 0$, so $^{\mathfrak{a}^{t}}I^{F}$ is also an ideal of $R$.
@@
    From part b), it follows that $^{\mathfrak{a}^{t}}I^{F}\subset{}^{\mathfrak{a}^{t}}({}^{\mathfrak{a}^{t}}I^{F})^{F}.$ Now, let $x\in{}^{\mathfrak{a}^{t}}({}^{\mathfrak{a}^{t}}I^{F})^{F}.$ Then for all $e\gg 0,$ $x^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset\mathfrak{a}^{\lceil tp^{e}\rceil}({}^{\mathfrak{a}^{t}}I^{F})^{[p^{e}]}\subset\mathfrak{a}^{\lceil tp^{e}\rceil}I^{[p^{e}]},$ where the last containment holds by checking the generators of $^{\mathfrak{a}^{t}}I^{F}$ and using that $^{\mathfrak{a}^{t}}I^{F}$ is a finitely generated ideal. ∎
@@
Unlike $\mathfrak{a}^{t}$-integral closure and $\mathfrak{a}^{t}$-tight closure, $\mathfrak{a}^{t}$-Frobenius closure cannot be computed using the corresponding closure ideal after going modulo the minimal primes. We use 2.6 to check that 3.13 may not hold true for $^{\mathfrak{a}^{t}}I^{F}$ and $I^{F\mathfrak{a}^{t}}$.
@@
Let $R=\mathbb{F}_{3}[X,Y,Z]/(Y^{6}-X^{3},X^{2}-Z^{2})$ and let $x,y,$ and $z$ denote the images of $X,Y,$ and $Z$ respectively in $R.$ Let $I=(x)$ and $\mathfrak{a}=(x,y^{2})$. We claim that $z\notin$ $^{\mathfrak{a}}I^{F}.$ Indeed, if $z\in$ $^{\mathfrak{a}}I^{F},$ then there exists $e$ such that $z^{3^{e}}(x,y^{2})^{3^{e}}\subset\mathfrak{a}^{3^{e}}I^{[3^{e}]}$. In particular, $z^{3^{e}}x^{3^{e}}\in\mathfrak{a}^{3^{e}}I^{[3^{e}]}=(x,y^{2})^{3^{e}}(x^{3^{e}})$. Therefore, $zx^{3^{e}-1}x^{3^{e}}=z(x^{2})^{(3^{e}-1)/2}x^{3^{e}}=z(z^{2})^{(3^{e}-1)/2}x^{3^{e}}=z^{3^{e}}x^{3^{e}}\in(x,y^{2})^{3^{e}}(x^{3^{e}})$. Since $x^{3^{e}}$ is a nonzerodivisor, we get $z\in(x,y^{2})^{3^{e}}:x^{3^{e}-1}=(x,y^{2})$ which is not true. Thus, the claim holds. Note that $\operatorname{Min}R=\{(x-z,y^{2}-z),(x+z,y^{2}+z)\}$ and so modulo either minimal prime, the image of $z$ is the same as the image of $x.$ Thus, $z+\mathfrak{p}\in((x)+\mathfrak{p})/\mathfrak{p}\subset$ $^{\mathfrak{a}}((x)+\mathfrak{p})^{F}/\mathfrak{p}$, but $z\not\in$ $^{\mathfrak{a}}(x)^{F}$.
Now let $I=(x+y^{2})$ and $\mathfrak{a}=(x,y^{2}).$ Then one can check that modulo any minimal prime $\mathfrak{p},$ the element $(1+\mathfrak{p})/\mathfrak{p}$ is in $((I+\mathfrak{p})/\mathfrak{p})^{F\mathfrak{a}}$ but $1\notin I^{F\mathfrak{a}}.$ In particular, it is enough to show that $(\mathfrak{a}^{3^{e}}+\mathfrak{p})/\mathfrak{p}\subset(I^{[3^{e}]}+\mathfrak{p})/\mathfrak{p}$ for every minimal prime $\mathfrak{p}$ and for all $e$ large but $\mathfrak{a}^{3^{e}}\not\subset I^{[3^{e}]}$ for any $e.$ Indeed, modulo any minimal prime $\mathfrak{p},$
@@
    If $t>\mu(\mathfrak{a})$ then
@@
    Thus, $(I^{F(f)})^{F(f)}=(I^{F}:(f)^{2})$ which is strictly larger than $I^{F(f)}=I^{F}:(f)$ under the given assumptions by Nakayama’s lemma, as $xf\in I^{F}$ implies $xf^{2}\in I^{F}\cdot(f)\subsetneq I^{F}$.
@@
4. Comparing the new closures
In this subsection, we compare the different adjusted closure operations with one another. All statements which involve the adjusted integral closure only are characteristic independent, but whenever the adjusted tight or Frobenius closure operations are mentioned, we tacitly assume the ambient ring is of prime characteristic $p>0$. Furthermore, in any statement where we invoke the $\mathfrak{a}^{t}$-integral closure, we assume $I$ and $\mathfrak{a}$ have positive height.
Remark 4.1.
By taking $c=1$ in the definition of the (adjusted) $\mathfrak{a}^{t}$-tight closures, we can easily see $I^{F\mathfrak{a}^{t}}\subset I^{*\mathfrak{a}^{t}}$ and $^{\mathfrak{a}^{t}}I^{F}\subset\,^{\mathfrak{a}^{t}}I^{*}$. Similarly, the adjusted $\mathfrak{a}^{t}$-tight closures are inside the corresponding adjusted $\mathfrak{a}^{t}$-integral closure, since if $x\in I^{*\mathfrak{a}^{t}}$, then for some $c\in R^{\circ}$ and all $e\gg 0$, $cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset I^{[p^{e}]}\subseteq I^{p^{e}}$ and so $x\in\overline{I}^{\mathfrak{a}^{t}}$ by Corollary 3.10, and similarly for $^{\mathfrak{a}^{t}}I^{*}\subset{}^{\mathfrak{a}^{t}}\overline{I}=\overline{I}$.
In general, we have the following inclusions of ideals, fixing $I$, $\mathfrak{a}$, and $t\in\mathbb{R}_{>0}$, and assuming $I$ and $\mathfrak{a}$ are of positive height.